The ratio of <1 to <2 is 4 to 5 respectively. Find the measure of <1. This is a 90 degree angle with angle 2 being smaller than angle 1.
I know the answer is 40 degrees. I have tried:
X +4x=90 but the answer doesn't match

Answer by jim_thompson5910(35256)

I'm assuming you mean "angle 1 and angle 2 add to 90 degrees"

so that would mean

4x+5x = 90
9x = 90
9x/9 = 90/9
x = 10

So angle 1 is 4x = 4*10 = 40 degrees